A spring-powered launcher has a spring with a spring constant of 150 N/cm. The launcher is prepared to launch a projectile by compressing the spring by 5.00 cm. If the mass of the projectile is 50.0 grams, then what is the velocity of the projectile when it is launched vertically from the launcher?
1
Expert's answer
2020-06-12T11:15:59-0400

Assume that energy is conserved:


"\frac{kx^2}{2}=\frac{mv^2}{2},\\\space\\
v=x\sqrt{\frac{k}{m}}=27\text{ m/s}."